OpenStudy (anonymous):

5r-3s=16 3r+5s=30 solve by elimination method.

OpenStudy (anonymous):

3(5r-3s= 16) = 15r-9s=48 [1] 5(3r+5s=30) = 15r+25s=150 [2] [2] - [1]gives 34s = 102 divide by 34 s = 3 Substitute 3 for s in 1 15r-27=48 (add 27 to both sides) 15r = 75 divide 15 on both sides r= 5 Answer s = 3, r = 5

OpenStudy (anonymous):

Thank you so much :)